OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 4) and (5, 1) B. (10, 6) and (1, 5) C. (10, 5) and (5, 1) D. (10, 6) and (5, 1)

OpenStudy (anonymous):

@Hero @ganeshie8 @Shay17

hero (hero):

Plug the points in to the given equation to see which pair work.

OpenStudy (anonymous):

more help? Its beeen 4 EVER since ive done these

OpenStudy (anonymous):

srry

hero (hero):

For example, for option A, first insert the point (x , y) = (10, 4) as follows: y = x - 4 4 = 10 - 4 4 = 6 False As soon as one of the points are false, the whole thing is false, so move on and test both points for point B

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

@Hero

hero (hero):

Can you show the work you did to arrive at B?